# DeRO: Dead Reckoning Based on Radar Odometry With Accelerometers Aided for Robot Localization

### v1

@@ IV-D Implementation
The detailed step-by-step practical implementation of the DeRO is outlined in Algorithm 1. In our approach, the procedure for each sensor is executed as soon as the corresponding sensor’s data becomes available.
@@
2:Performing course alignment algorithm.
@@
5:         Compensating biases $\hat{\omega}^{b}=\bar{\omega}^{b}-\hat{b}_{g}$.
@@
15:              Registration $\rightarrow\hat{p}^{r_{k}}_{r_{k-M}},\hat{\sigma}_{p_{r}}^{2}$.
16:              Radar measurement update with $\breve{H}_{r}$ and $\breve{P}_{k}^{-}$.
17:              Reset $scanCount$.
18:         end if
19:     end if
20:     if Accelerometer is available then
21:         Compensating linear acceleration $\leftarrow$ (41).
22:         Calculating tilt angles $\rightarrow\hat{\phi}_{a},\hat{\theta}_{a}$.
23:         Accelerometer measurement update $\leftarrow$ $H_{a},P_{k}^{-}$.
24:     end if
25:end for
@@ V-A Open-source datasets
To verify the proposed DeRO algorithm, five open source datasets provided by [23] was used, namely Carried 1 to 5. Here, we briefly describe overview of the datasets; for more detailed information about the setup, readers are encouraged to refer to the mentioned paper. These datasets were recorded in an office building using a hand-held sensor platform equipped with a 4D milimeter-wave FMCW radar (TI IWR6843AOP) and an IMU (Analog Devices ADIS16448). The IMU provides data at approximately 400 Hz, while the radar captures 4D point clouds at 10 Hz. In addition, the radar has a FOV of 120 deg for both azimuth and elevation angles. Furthermore, the sensor platform includes a monocular camera, which is utilized to generate a pseudo ground truth via the visual inertial navigation system (VINS) algorithm with loop closure. It is worth noting that although the IMU also provides barometric measurements, we did not utilize them.
[…]
@@ V-B1 Evaluation of 2D Trajectory
Figure 3 illustrates the top-view of the aligned X-Y trajectory estimates obtained by the EKF RIO and SC-IEKF DeRO. Overall, Both methods demonstrate decent estimation performance. However, our proposed method exhibits superior performance, particularly evident when the sensor platform traces a straight loop from $(x,y)=(7,0)$ to $(x,y)=(7,-25)$. Furthermore, as the start and end points of the trajectory coincide, we can relatively judge the filter’s accuracy solely by examining the final estimated positions. Specifically, SC-IEKF yields an approximate position of $(\hat{x},\hat{y},\hat{z})=(-0.82,-0.84,0.23)$, while EKF RIO yields approximately $(5.49,-1.42,1.71)$, corresponding to distance errors of approximately 1.2 m and 5.9 m, respectively. Thus, our proposed method reduces the final position distance error from the conventional approach by approximately 79.6%.
@@ V-B2 Evaluation of Relative Errors
The relative errors in terms of translation and rotation are depicted in Fig. 4 using boxplots. Our method consistently outperforms the EKF RIO in both categories. Although the translation error boxplot shows a slightly larger error and uncertainty for the SC-IEKF DeRO after the sensor platform travels 50 m compared to its competitor, the EKF RIO exhibits rapid position drift at distances of {100, 150, 200} m, whereas our proposed method maintains an acceptable error level from start to finish. This trend is also evident in the rotation error plot.
@@
Dataset SC-IEKF DeRO EKF RIO [11]
@@
Carried 1 (287 m, 273 s) $\boldsymbol{0.722}$ $\boldsymbol{3.547}$ $1.585$ $8.653$
Carried 2 (451 m, 392 s) $\boldsymbol{1.279}$ $\boldsymbol{6.925}$ $3.855$ $21.049$
Carried 3 (235 m, 171 s) $\boldsymbol{0.914}$ $\boldsymbol{4.162}$ $1.256$ $6.130$
Carried 4 (311 m, 220 s) $\boldsymbol{1.128}$ $\boldsymbol{4.653}$ $1.696$ $8.335$
Carried 5 (228 m, 172 s) $\boldsymbol{1.228}$ $\boldsymbol{4.544}$ $1.511$ $5.535$
Mean ATE $\boldsymbol{1.05}$ $\boldsymbol{4.76}$ $1.98$ $9.94$
@@ V-B3 Evaluation of ATE
Table I summarizes the ATE in terms of both translation and rotation for the two considered algorithms across all datasets. Once again, the SC-IEKF DeRO completely outperforms the EKF RIO in all trials, especially in the Carried 1 and 2 datasets, which features numerous turns. For instance, in the Carried 2 field test (the longest trajectory with 451 m distance and 392 seconds duration), our developed method yields a translation error of 1.279 m and rotation error of 6.925 deg, compared to 3.855 m and 21.049 deg for the EKF RIO. This represents a translation error reduction of 66.8% and a rotation error reduction of 67.1%. Overall, across the mean ATE of the five experiments, our DeRO approach reduces the translation error by approximately 47% and the rotation error by 52%.

### v2

@@ IV-D Implementation
The detailed step-by-step practical implementation of the DeRO is outlined in Algorithm 1. Here, we assume accelerometer data is collected faster than radar data, allowing approximate timestamp synchronization. This enables both data sources to be used concurrently for KF’s measurement update. Furthermore, to enhance the performance of ICP algorithm, we utilize the estimations from the previous $M$ steps to calculate the initial guessed pose.
@@
2:Performing coarse alignment algorithm $\rightarrow\hat{b}_{g},\hat{b}_{a}$.
@@
5:         Compensating gyroscope biases $\hat{\omega}^{b}=\bar{\omega}^{b}-\hat{b}_{g}$.
@@
15:              Registration $\rightarrow\hat{p}_{r_{k}}^{r_{k-M}},\hat{\sigma}_{p_{r}}^{2}$.
16:              Computing $\breve{H}_{r}$.
17:              if Accelerometer is available then
18:                  Compensating accelerometer bias and
19:      linear acceleration $\leftarrow$ (26).
20:                  Calculating $\rightarrow\hat{\phi}_{a},\hat{\theta}_{a}$ and $H_{a}$.
21:                  Stacking $H_{a}$ and $\breve{H}_{r}$.
22:              end if
23:              Estimation update with stacked measurement
24:      matrix and $\breve{P}_{k}^{-}$.
25:              Reset $scanCount$.
26:         end if
27:     end if
28:end for
@@ V-A Open-source datasets
To verify the proposed DeRO algorithm, five open source datasets provided by [23] was used, namely Carried 1 to 5. We briefly describe overview of the datasets; for more detailed information about the setup, readers are encouraged to refer to the mentioned paper. These datasets were recorded in an office building using a hand-held sensor platform equipped with a 4D milimeter-wave FMCW radar (TI IWR6843AOP) and an IMU (Analog Devices ADIS16448). The IMU provides data at approximately 400 Hz, while the radar captures 4D point clouds at 10 Hz. In addition, the radar has a field of view of 120 deg for both azimuth and elevation angles. Furthermore, the sensor platform includes a monocular camera, which is utilized to generate a pseudo ground truth via the visual inertial navigation system (VINS) algorithm with loop closure.
We implemented DeRO in ROS 2 Foxy [14] using C++. The datasets were processed on a Linux PC equipped with an Intel i9-12900K CPU running at 3.20 GHz. Additionally, to ensure fairness in comparison, we independently replicated the EKF RIO instead of relying on the provided code by the authors. Note that we maintained consistency with the algorithm’s parameters, such as RANSAC/LSQ, sensor noise density, initialization values, etc., as specified in [11]. For scan matching, we employed a standard iterative closest point (ICP) algorithm with the window size $M=$ 3. The rationale behind this number is that the radar measurement frequency in the public dataset is approximately 10 Hz. We aim for the filter to update the estimation twice per second. Based on the quality metric provided by the ICP output, we manually tuned the measurement noise variance $\sigma_{p_{r}}^{2}$. Besides, we selected $\gamma=$ 0.059 m/s2, as recommended by [13], for the noise covariance adaptive rule.
@@ V-B1 Evaluation of 2D Trajectory
Figure 3 illustrates the top-view of the aligned X-Y trajectory estimates obtained by the EKF RIO and SC-IEKF DeRO. Overall, Both methods demonstrate decent estimation performance. However, our proposed method exhibits superior performance, particularly evident when the sensor platform traces a straight loop from $(x,y)=(7,0)$ to $(x,y)=(7,-25)$. Furthermore, as the start and end points of the trajectory coincide, we can relatively judge the filter’s accuracy solely by examining the final estimated positions. Specifically, SC-IEKF DeRO yields an approximate position of $(\hat{x},\hat{y},\hat{z})=(-1.07,-0.29,0.75)$, while EKF RIO yields approximately $(3.15,-1.37,1.34)$, corresponding to distance errors of approximately 1.3 m and 3.7 m, respectively. Thus, our proposed method reduces the final position distance error from the conventional approach by approximately 65%.
@@ V-B2 Evaluation of Relative Errors
The relative errors in terms of translation and rotation are depicted in Fig. 4 using boxplots. Our method consistently outperforms the EKF RIO in both categories. Notably, although the translation error of SC-IEKF DeRO shows slightly higher error and uncertainty after the sensor platform travels 50 m, it reliably maintains the error within 2.5 m thereafter. In contrast, the EKF RIO experiences a rapid position drift, reaching up to 4 m over a distances of 200 m. For rotation estimation, while the proposed algorithm demonstrates a smaller mean, it exhibits a larger interquartile range due to the slower update rate of the ICP compared to the traditional method.
@@
Dataset EKF RIO [11] SC-IEKF DeRO
@@
Carried 1 (287 m, 273 s) $1.075$ $5.370$ $\boldsymbol{0.617}$ $\boldsymbol{2.734}$
Carried 2 (451 m, 392 s) $2.724$ $14.768$ $\boldsymbol{0.709}$ $\boldsymbol{2.869}$
Carried 3 (235 m, 171 s) $1.110$ $5.741$ $\boldsymbol{0.464}$ $\boldsymbol{1.853}$
Carried 4 (311 m, 220 s) $1.547$ $7.448$ $\boldsymbol{0.491}$ $\boldsymbol{2.659}$
Carried 5 (228 m, 172 s) $1.495$ $5.331$ $\boldsymbol{0.741}$ $\boldsymbol{3.093}$
Mean ATE $1.590$ $7.731$ $\boldsymbol{0.604}$ $\boldsymbol{2.641}$
@@ V-B3 Evaluation of ATE
Table I summarizes the ATE in terms of both translation and rotation for the two considered algorithms across all datasets. Once again, the SC-IEKF DeRO completely outperforms the EKF RIO in all trials, especially in the Carried 1 and 2 datasets, which features numerous turns. For instance, in the Carried 2 field test (the longest trajectory with 451 m distance and 392 seconds duration), our developed method yields a translation error of 0.709 m and rotation error of 2.869 deg, compared to 2.724 m and 14.768 deg for the EKF RIO. This represents a translation error reduction of 74% and a rotation error reduction of 80%. Overall, across the mean ATE of the five experiments, our DeRO approach reduces the translation error by approximately 62% and the rotation error by 66%.